### v1

# The heat flow on glued spaces with varying dimension

Abstract
We examine under which conditions the canonical heat flow on glued manifolds is ergodic and irreducible. Glued manifolds are spaces consisting of manifolds of varying dimension connected by a weakly doubling measure. Moreover, we construct a non-local perimeter functional, the heat excess, to raise the question of its $\Gamma-$convergence to the standard perimeter functional.
In this context, we connect our work to the previous work on the convergence of perimeter functionals, approximations, and existence of heat kernels, as well as short-time expansions of Brownian motion.
Keywords: Heat flow, glued manifolds, stratification, ergodic flows, Brownian motion, manifold learning.
MSC2020: 58J35 (Primary); 58A35; 37A30 (Secondary)
1 Motivation
The heat equation is one of the most important and most often applied partial differential equations. This equation has applications in a vast amount of areas and fields. In this note, we analyze the existence of an ergodic heat flow on so called weighted glued manifolds. These include spaces which consist of manifolds with different dimensions that are connected through weights. The idea behind these spaces is to incorporate a general setting for e.g. manifold learning and similar applications where data is distributed locally in the shape of manifolds but can vary globally. The weights make sure that each component distributes in the same way and that the data is overall continuous in some sense. To state our main theorem, Theorem 2.2, in an informal way, we will prove
Theorem.
Let $(X,d,\mu)$ be the union of weighted smooth manifolds with transversal intersections s.t. $\mu$ fulfills a weak doubling condition. Then, the canonical heat flow on $X$ is ergodic and irreducible.
Since the heat flow is ergodic and irreducible it is possible to investigate the short-term behavior of the heat flow and investigate the $\Gamma-$ (or Mosco-)convergence of a non-local perimeter functional to the local perimeter. With this in hand, one can construct a classifier on these spaces. Moreover, the definitions carry over to a more general setting of metric measure spaces instead of smooth compact manifolds. For a more details, see 5.
1.1 Structure of the note
First, in Section 2, the setting in which we are working and the basic definitions are stated together with our main theorem in a precise form, Theorem 2.2. We consider a class of glued manifolds with mild assumptions on their weights to ensure connectedness in a weak sense.
Additionally, we recapitulate the standard theorems and properties which we apply later. The advanced theorems and notations are introduced in their respective sections as to not overload Section 2.
Afterwards, in Section 3, we prove our main theorem in a simple setting in three steps. We begin with the definition of the heat flow and associated operators and spaces. Next, we prove continuity in a weak sense via capacity bounds and conclude by a compact embedding in the sense of Rellich-Kondrachov. This embedding yields the ergodicity and irreducibility.
Roughly, the idea is to use the local structure of the manifolds to show that only constants remain static under the flow. Together with spectral theory of the resolvent associated with the flow, this allows to prove a spectral gap and exponential convergence to constants for every initial condition. This convergence, in turn, implies ergodicity and irreducibility.
In Section 4, we point out ways to extend the results to more general weights for glued spaces and prove that our procedure from Section 3 can be iterated to the setting of multiple glued manifolds as long as they only intersect pairwise.
Next, in Section 5, we present the definitions and basic theory needed to state a conjecture about the $\Gamma-$convergence (or Mosco-convergence) of a non-local perimeter functional to the perimeter functional in glued spaces and general metric measure spaces. This functional is known as the heat excess and uses the heat flow constructed beforehand.
Finally, in Section 6, we discuss related work in the different areas of this note and connections to the different fields of study where these problems occur.
2 Setting and main theorem
We will now give the setting for the metric measure space $(X,d,\mu)$.
Our domain $X$ is the union of two embedded Euclidean manifolds that can vary in dimension and intersect. To be precise,
$X:=M_{1}\cup M_{2}$
where $M_{i}$ are compact, connected, oriented manifolds with smooth boundary. The dimension of $M_{i}$ is $n_{i}$ and the intersection $\emptyset\neq L:=M_{1}\cap M_{2}$ lies in the inner of $M_{i}$ and is a smooth manifold of dimension $k$. This is the case if e.g. the intersection is transversal in the surrounding space or a smooth manifold containing an open neighborhood of the intersection, c.f. [29, Chapter 5].
The (intrinsic) distance $d$ is defined as the glued distance which is obtained by taking the infimum over the length of all connecting paths between points. All balls $B$ in this note are w.r.t. the intrinsic distance $d$ on $X$.
Lastly, $\mu$ takes the role of a volume measure on $X$ that connects the manifolds and is composed of the weighted volume measures
$\mu:=\omega_{1}\cdot\text{vol}_{M_{1}}+\omega_{2}\cdot\text{vol}_{M_{2}}$
where the weights $\omega_{i}\in A_{2}(M_{i}),\omega_{i}>0$ are Muckenhoupt, see Definition 2.9. Additionally, we require that $\mu(L)=0$.
Moreover, we assume that the measure $\mu$ satisfies a weak continuity in the style of a doubling condition:
Definition 2.1 ($N$-doubling condition).
A measure $\mu$ on a metric measure space $(X,d,\mu)$ is called $N-$doubling if for all $x\in X$ and radii $r>0$ the metric balls are doubling up to a function $N(r):$
$\mu(B_{3r}(x))\leq N(2r)\mu(B_{r}(x)).$
This ensures a weak type of continuity of $\mu$.
Remark.
The $N$-doubling condition can be formulated with the Euclidean (extrinsic) distance. In this case one would need to strengthen the transversality into a metric-quantitative version, i.e. there exist balls near the intersection which do not intersect other manifolds.
We note that $\mu$ is a positive Radon measure with $\operatorname*{supp}\mu=X$ and that $L^{2}(X,\mu)$ is a Hilbert space.
Remark.
Alternatively, one could consider another connection of the manifolds instead of the weights. It is possible to get a continuity over the intersection by collapsing an area around it into a lower dimensional submanifold and changing the distances and topology accordingly. In this note, we decided for the approach with weights as it satisfies a continuity over the whole domain and fits more naturally into the framework of manifold learning.
Now, we are able to state the main theorem of this note.
Theorem 2.2 (Heat flow on the glued manifold).
Let $X=M_{1}\cup M_{2}$ with the intrinsic distance $d$ be a given transversal union of two smooth compact manifolds with boundary s.t. the intersection $L:=M_{1}\cap M_{2}$ lies in the inner of both of them. Moreover, assume non-negative weights $\omega_{1}\in A_{2}(M_{1})$ and $\omega_{2}\in A_{2}(M_{2}),\omega_{i}>0$ s.t. $\omega_{2}$ is bounded from above and below with $k=n_{2}-1$ where w.l.o.g. $n_{1}\geq n_{2}$. Additionally, assume that
$\mu:=\omega_{1}\cdot\text{vol}_{M_{1}}+\omega_{2}\cdot\text{vol}_{M_{2}}$
fulfills $\mu(L)=0$ and is $N-$doubling with $N\in L^{1}_{loc}(\mathbb{R}_{\geq 0})$.
Then, the canonical heat flow is ergodic and irreducible.
The construction of the heat flow is done via its semigroup which itself stems from the natural Dirichlet form $\mathcal{E}$ on $X$:
$\operatorname*{\mathcal{E}}(f):=\int\limits_{X}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu:=\int\limits_{M_{1}}|\nabla^{M_{1}}f|^{2}\mathop{}\!\mathrm{d}\mu_{1}+\int\limits_{M_{2}}|\nabla^{M_{2}}f|^{2}\mathop{}\!\mathrm{d}\mu_{2}.$
This is to say that $\operatorname*{\mathcal{E}}$ is the manifold gradient on each manifold $M_{i}$ combined with the partial measure $\mu_{i}:=\omega_{i}\cdot\text{vol}_{M_{i}}$. For a general definition of Dirichlet forms and their properties see Definition 2.4 and following.
[…]
Figure 1: Example of a glued manifold
Example 2.3.
A typical example would be a one-dimensional line intersecting a two-dimensional disk with the density $\frac{1}{|x|}$ that lifts the dimensions at their intersection. This is so that their dimensions don’t differ too much and the measure is continuous in some way. A visualization can be seen in Figure 1. We will see that if the difference in dimension at their intersection is less than one, the heat flow connects both of them and is ergodic. In the case that the difference in their dimension is too large, the capacity of the intersection w.r.t. one of both manifolds is zero and thus the heat flow does not need to be ergodic.
Now, we will give the precise definition of Dirichlet forms as well as of Muckenhoupt weights. The following definitions can be compared to [20, 60, 62].
Definition 2.4 (Dirichlet form).
A symmetric form (non-negative definite densely defined symmetric bilinear form) $(\mathcal{E},D(\mathcal{E}))$ on $L^{2}$ is called a Dirichlet form if it is closed and Markovian, i.e.
$D(\mathcal{E})\text{ with norm }\mathcal{E}_{1}(u):=\left\lVert u\right\rVert_{L^{2}}+\mathcal{E}(u,u)\text{ is complete}$
and for
$u\in D(\mathcal{E}),v=(0\lor u)\land 1,\text{ we have }v\in D(\mathcal{E}),\mathcal{E}(v,v)\leq\mathcal{E}(u,u).$
A core $\mathcal{C}$ of $\mathcal{E}$ is a dense subset of $D(\mathcal{E})\cap C^{0}$ (in $D(\mathcal{E})$ w.r.t. $\mathcal{E}_{1}$ and in $C^{0}$ w.r.t. $L^{\infty}-$norm) and we call $\mathcal{E}$ regular if a core exists.
Lastly, a Dirichlet form $\mathcal{E}$ is called strongly local if for $u,v\in D(\mathcal{E})$ compactly supported and $u$ constant in a neighborhood of $v$: $\mathcal{E}(u,v)=0$. Furthermore, it is called strictly local if the associated energy metric $d_{\mathcal{E}}$ (see below) is a metric and induces the same topology as $d$.
We will write $\mathcal{E}(u)$ for $\mathcal{E}(u,u)$. For closed forms, the Markovian condition can be rephrased as $\mathcal{E}$ being non-increasing under normal contractions.
Now, we will define the energy metric and what we mean when we call a function (locally) Lipschitz. On the core and thus by extension on the domain, we define the energy measure $\Gamma(u)$ which is a Radon measure, see [60, 62], via:
$\int_{X}\varphi\mathop{}\!\mathrm{d}\Gamma(u):=\mathcal{E}(\varphi u,u)-\frac{1}{2}\mathcal{E}(u^{2},\varphi).$
This can also be seen as a carré-du-champ-operator if ones looks at the bilinear form obtained using a polarization formula and for $u,v\in D(\mathcal{E})$
$\int_{X}\varphi\mathop{}\!\mathrm{d}\Gamma(u,v)=\frac{1}{2}\left(\mathcal{E}(\varphi u,v)+\mathcal{E}(\varphi v,u)-\frac{1}{2}\mathcal{E}(\varphi,uv)\right).$
Given $\Gamma$, we now can define the energy metric which sometimes is also called intrinsic metric for a regular Dirichlet form:
$d_{\mathcal{E}}(x,y):=\sup\left\{u(x)-u(y)\left|u\in\mathcal{C},\frac{\mathop{}\!\mathrm{d}\Gamma(u)}{\mathop{}\!\mathrm{d}\mu}\leq 1\right.\right\}.$
This notion means that $\Gamma(u)\ll\mu$ and that the Radon-Nikodym derivative is bounded by $1$. A priori, this energy metric does not have to be a metric as it can be infinite or zero for different $x,y$.
We note that $\mathcal{E}$ in the form above for the setting of manifolds is in fact a Dirichlet form and is regular, strongly local and strictly local. In particular, we could also use the energetic metric instead of the glued distance.
Finally, we call a function $u\in D(\mathcal{E})$ Lipschitz, $u\in\operatorname*{Lip}$, if $\Gamma(u)\ll\mu$ and the Radon-Nikodym derivative is bounded. This bound is then called the Lipschitz constant. In this case, we write $|\nabla u|^{2}:=\frac{\mathop{}\!\mathrm{d}\Gamma(u)}{\mathop{}\!\mathrm{d}\mu}$ and $|\nabla u|:=\sqrt{|\nabla u|^{2}}$. This coincides with our intuition that $\mathcal{E}$ is the $H^{1}-$seminorm and $\Gamma$ the square of the derivative.
For the connectedness of $(X,d,\mu)$ we will need a Poincare inequality.
Definition 2.5 ($2$-Poincare inequality).
We say that $X$ satisfies a $2$-Poincare inequality if for each $f\in\operatorname*{Lip}$ there exists $C>0$ and a dilation factor $\lambda\geq 1$ s.t. for each ball $B\subset X$:
$\fint_{B}|f-f_{B}|^{2}\mathop{}\!\mathrm{d}\mu$ $\leq Cr^{2}\fint_{\lambda B}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu.$
Remark.
For generalizations of the definitions, see Section 5.
Note, if $X$ satisfies a Poincare inequality it has to be connected up to negligible sets. Assume it would not be connected up to a null set and let $X_{1}$ be one component with $f=\chi_{X_{1}}$ and $|\nabla f|=0$. Then, this would be a contradiction as $\mu(X_{1})>0$ and $01>f_{B}>$ 0 for $B$ large enough.
If not specified we will assume the dilation factor $\lambda$ to be one.
The heat flow is defined using its semigroup which originates in the Dirichlet form. In summary, this can be done as follows. Given the Dirichlet form $\operatorname*{\mathcal{E}}$, we define the Laplace-Beltrami operator $\operatorname*{\mathcal{A}}$ with its domain $D(\operatorname*{\mathcal{A}})$ in the following way
$D(\operatorname*{\mathcal{A}}):=\{f\in D(\operatorname*{\mathcal{E}})|\exists\operatorname*{\mathcal{A}}f\in L^{2}:(\operatorname*{\mathcal{A}}f,g)=-\operatorname*{\mathcal{E}}(f,g)\;\forall g\in D(\operatorname*{\mathcal{E}})\}.$
By definition, this is a selfadjoint symmetric non-positive operator which induces an analytic semigroup $S_{h}$ which we call the heat semigroup.
In Section 3, we will also show continuity and ergodicity properties of these objects. For this, we will need capacity bounds.
Definition 2.6 (Capacities).
The relative $2-\mu-$capacity is defined on compact sets and extended to measurable sets in the usual way:
$\text{Cap}_{2,\mu}(K,\Omega):=\inf\left\{\left.\int\limits_{\Omega}|\nabla u|^{2}\mathop{}\!\mathrm{d}\mu\right|u\in\operatorname*{Lip}\cap C_{c}(\Omega),u\geq\mathds{1}_{K}\right\}.$
For an overview of the theory of capacities and more general definitions, see [36, 40, 41].
Additionally, we want to show that our constructed heat flow connects the whole space. The right notion to make this precise is the notion of ergodicity and irreducibility, see [20]. In the following, let $S_{h}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ be a semigroup.
Definition 2.7 (Irreducibility).
We call a measurable set $E\subset X$ $S_{h}$ invariant if for all $f\in L^{2},h>0$ we have
[…]
Then, $S_{h}$ is irreducible if for all $S_{h}$ invariant sets $E$ we have $\mu(E)=0$ or $\mu(E^{c})=0$.
Definition 2.8 (Ergodicity).
We call $S_{h}$ ergodic if the set function $T_{t}$ defined by
$T_{t}:X\to X,\quad E\mapsto\operatorname*{supp}(S_{t}\mathds{1}_{E})$
is ergodic for each fixed $t>0$, i.e. if for a measurable set $E$ with $T_{t}^{-1}(E)=E$ we have $\mu(E)=0$ or $\mu(E^{c})=0$.
Lastly, for our setting of glued manifolds with weights, we need to define what is means to be a Muckenhoupt weight.
Definition 2.9 (Muckenhoupt weights).
We say $\omega$ is in the class $A_{2}(M)$ of $2-$Muckenhoupt weights on a smooth compact manifold $M$ if for all balls $B$
$\fint\limits_{B}\omega_{i}\mathop{}\!\mathrm{d}\text{vol}_{M_{i}}\cdot\fint\limits_{B}\frac{1}{\omega_{i}}\mathop{}\!\mathrm{d}\text{vol}_{M_{i}}\lesssim 1.$
Remark.
For the typical weight $\omega=|x|^{\alpha}$ in $\mathbb{R}^{d}$, we have $\omega\in A_{2}$ iff $\alpha\in(-n,n)$.
2.1 Basic theorems for $A_{2}-$weights on manifolds
In this section, we will show that, if we have a smooth compact manifold $M$ and a Muckenhoupt weight $\omega\in A_{2}(M)$, then the measure $\mu=\omega\cdot\text{vol}_{M}$ is doubling and fulfills the $2-$Poincare inequality and a reverse Hölder-inequality. These theorems are mostly due to [36, Chapter 15] where they are carried out in the case of a flat Euclidean space.
Lemma 2.10 ($A_{2}$ doubling).
Muckenhoupt weights satisfy a volume doubling condition.
Proof.
For $E\subset B$ where $B$ is a ball we have:
$\text{vol}(E)$ $=\int\limits_{E}\frac{\omega^{\frac{1}{2}}}{\omega^{\frac{1}{2}}}\mathop{}\!\mathrm{d}x$
$\leq\sqrt{\mu(E)}\sqrt{\int\limits_{E}\frac{1}{\omega}\mathop{}\!\mathrm{d}x}$
$\leq\sqrt{\mu(E)\text{vol}(B)}\sqrt{\fint\limits_{B}\frac{1}{\omega}\mathop{}\!\mathrm{d}x}$
$\lesssim\sqrt{\mu(E)}\text{vol}(B)\sqrt{\frac{1}{\mu(B)}}$
$\Rightarrow\mu(B)$ $\lesssim\left(\frac{\text{vol}(B)}{\text{vol}(E)}\right)^{2}\mu(E).$
A direct consequence is the volume doubling for $E=B_{r},B=B_{2r}$ as the volume measure of a compact manifold is doubling:
$\mu(B_{2r})\lesssim\mu(B_{r}).$
The other properties can be proven locally on charts as we are only interested in inequalities on balls with a radius bounded above by an arbitrary but fixed bound.
The balls $B_{r}^{M}$ on $M$ are transformed by transition map $\phi$ to $A_{r}$. Since $\phi$ can be chosen Lipschitz, there exists $C>1$ s.t. $B_{\frac{r}{C}}\subset A_{r}\subset B_{Cr}$ and thus (since the metric is bounded)
$\text{vol}(A_{r})\simeq|A_{r}|\simeq|B_{r}|.$
Therefore,
$\int\limits_{B_{r}}\omega\mathop{}\!\mathrm{d}x\int\limits_{B_{r}}\frac{1}{\omega}\mathop{}\!\mathrm{d}x$ $\leq\int\limits_{A_{Cr}}\omega\mathop{}\!\mathrm{d}x\int\limits_{A_{Cr}}\frac{1}{\omega}\mathop{}\!\mathrm{d}x$
$\lesssim|A_{Cr}|^{2}\lesssim|B_{r}|^{2}.$
We conclude that $\omega$ (to be concrete $\omega\circ\phi^{-1}$) is also a Muckenhoupt weight on the flat chart, similar for $\omega\in A_{p}$, see Section 5. Hence, we can follow the usual proofs of [36, Chapter 15] to get a $p-$Poincare inequality without dilation. First, we can prove the doubling condition and also the reverse doubling condition which states for each ball $B$ and measurable subset $E\subseteq B$ there exists $10<q<$ 1 s.t.
$\frac{\mu(E)}{\mu(B)}\leq\left(\frac{|E|}{|B|}\right)^{q}.$
From this, we directly conclude the reverse Hölder inequality: There is $r>1$ s.t.
$\left(\fint_{B}\omega^{r}\mathop{}\!\mathrm{d}x\right)^{\frac{1}{r}}\lesssim\fint_{B}\omega\mathop{}\!\mathrm{d}x.$
This inequality can be used to obtain the open end property, i.e. if $\omega\in A_{p}$ with $p>1$ then $\omega\in A_{p-\varepsilon}$ for some $\varepsilon>0$. Concluding, this yields an equivalent characterization via the continuity of the Hardi-Littlewood maximal function on $L^{p}(\mu)-$functions, which finally proves a Poincare inequality on local balls for $\mu=\omega\cdot\text{vol}_{M}$ with $\omega\in A_{2}$:
$\forall r<R:\int\limits_{B_{r}}|f-(f)_{B_{r}}|^{2}\mathop{}\!\mathrm{d}\mu\lesssim r^{2}\int\limits_{B_{r}}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu.$
3 Proof of main theorem
3.1 Definitions
Given the setting of Section 2, we will prove our main theorem, Theorem 2.2. This theorem can be extended into the more general case, for this see Section 4.
The proof consists of three different steps to prove properties of the resulting space as well as of the operators and semi group on it. This will be done in the following. First, we define the relevant operators and function spaces. Afterwards, we will prove a continuity in a capacity sense. This, we will use to show a Rellich-Kondrachov embedding which yields ergodicity and irreducibility.
We start with the definition of the heat flow via a given Dirichlet form using the Laplace-Beltrami operator.
We define the energy norm
$\left\lVert f\right\rVert_{H^{1}_{*}}:=\sqrt{\left\lVert f\right\rVert_{L^{2}(X,\mu)}^{2}+\operatorname*{\mathcal{E}}(f)}$
on the domain of $\mathcal{E}$
$D(\operatorname*{\mathcal{E}}):=\overline{C^{\infty}(\mathbb{R}^{d})}^{\left\lVert\cdot\right\rVert_{H^{1}_{*}}}=:H^{1}_{*}.$
Here, $d$ is the dimension of the surrounding space where $X$ is embedded in. The asterisk symbolizes the continuity which we will see in Section 3.2. Heuristically, this can be seen as a pointwise continuity on the intersection and Neumann conditions at the boundary as well as flow conditions at points where the manifolds meet.
Additionally, we define the Laplace-Beltrami operator $\operatorname*{\mathcal{A}}$ with its domain $D(\operatorname*{\mathcal{A}})$ via the Dirichlet form in the following way
$D(\operatorname*{\mathcal{A}}):=\{f\in D(\operatorname*{\mathcal{E}})|\exists\operatorname*{\mathcal{A}}f\in L^{2}(X,\mu):(\operatorname*{\mathcal{A}}f,g)=-\operatorname*{\mathcal{E}}(f,g)\;\forall g\in D(\operatorname*{\mathcal{E}})\}.$
By definition, this is a selfadjoint symmetric non-positive operator which induces an analytic semigroup $S_{h}$ that we call the heat semigroup. The properties of $\operatorname*{\mathcal{A}}$ are computed via the resolvent
$R_{\operatorname*{\mathcal{A}}}=(\operatorname*{Id}-\operatorname*{\mathcal{A}})^{-1}:L^{2}(X,\mu)\to D(\operatorname*{\mathcal{A}})\subseteq D(\operatorname*{\mathcal{E}})$
which satisfies $(f,g)=(R_{\operatorname*{\mathcal{A}}}f,g)+\operatorname*{\mathcal{E}}(R_{A}f,g),Af=f-R_{\operatorname*{\mathcal{A}}}^{-1}f$ for $f\in D(\operatorname*{\mathcal{A}})$.
In Section 3.3, we will show and use that $R_{\operatorname*{\mathcal{A}}}$ is compact to prove a spectral gap on $\operatorname*{\mathcal{A}}$.
3.2 Capacity bounds
In this part, we show that the $N$-doubling condition implies a capacity equivalence if $N$ is integrable:
Lemma 3.1 (Capacity equivalence).
Under the assumptions of Theorem 2.2, the capacity of the intersection is positive in both manifolds:
$\text{Cap}_{2,\mu_{1}}(L,X)>0,\quad\text{Cap}_{2,\mu_{2}}(L,X)>0.$
Proof.
We recall that intersection of the manifolds $L$ is a smooth manifold of dimension $k$.
In the following, we treat the intersection as a linear subspace $L\subseteq\mathbb{R}^{k}$. This is rectified by considering charts of the manifolds s.t. they agree on the intersection $L$ and possess smooth (bi-Lipschitz) transition maps which map $L$ into the linear subspace. By compactness of $L$ and $M_{i}$, finitely many suffice and the capacity of $L$ in $M_{i}$ is finite iff there is one chart on which the relative capacity is finite.
Moreover, by continuity of the metric, the charts are chosen s.t. $\sqrt{\det g}$ is uniformly bounded from above and below and s.t. they agree on $L$ for both manifolds. Thus, we will work with the Lebesgue measure $\mathcal{L}^{n}$ instead of the volume measure $\text{vol}_{M}$ in this section as the calculations are equivalent up to a factor.
First, we will need an equivalence of the partial measures close to the intersection.
Let $R^{*}(x)$ be small enough s.t. for every $x^{*}\in L,R<R^{*}$ there is a point $x\in\partial B_{\frac{R}{2}}(x^{*})\cap M_{1}$ with $B_{\frac{R}{2}}(x)\subset M_{1}$. Then, we can estimate the partial measure $\mu_{i}$ on balls on the intersection. W.l.o.g. only written down in the direction $M_{2}\leq M_{1}$:
$\mu_{2}(B_{R}(x^{*}))$ $\leq\mu(B_{R}(x^{*}))\leq\mu(B_{\frac{3}{2}R}(x))\leq N(R)\mu(B_{\frac{R}{2}}(x)))$
$=N(R)\mu_{1}(B_{\frac{R}{2}}(x))\leq N(R)\mu_{1}(B_{R}(x^{*})).$
This means that we have
$\forall x\in L,R<R^{*}(x):N(R)^{-1}\leq\frac{\mu_{1}(B_{R}(x))}{\mu_{2}(B_{R}(x))}\leq N(R).$
Remark.
$R^{*}(x)$ can be bounded uniformly from below since $M_{i}$ and thus also $L$ are compact manifolds.
The point $x$ needs to exist by dimensional arguments as $k+1=n_{2}<n_{1}$ and thus the dimension $d^{\prime}$ of the manifold in which $M_{1}$ and $M_{2}$ are transversal is larger than $n_{1}$ and $n_{2}$ as $d^{\prime}=n_{1}+n_{2}-k$. Therefore, there exists a direction solely in $T_{x^{*}}M_{1}$ for each $x^{*}\in L$ which yields existence of $x$ by smoothness of the manifolds.
To compute the capacity, it proves useful to consider a thickened version of the intersection $L_{R}:=\bigcup\limits_{x\in L}B_{R}(x)$ as we will estimate the relative capacity $\text{Cap}_{2,\mu}(L_{r},L_{R})$ on $L_{R}$ via an approximation $L_{r}\to L$ as $r\searrow 0$. In the following, we will estimate the capacity from below on both manifolds in order to ensure that the manifolds are sufficiently connected. The techniques are based on [36, Chapter 2] where the case of balls is proven.
For this, we need a representation formula.
In this, we will work on $\mathbb{R}^{n}$ by the estimate of the volume measure and the smoothness of the transition map this is justified as the calculations carry over to the charts and manifolds. Take $u\in C^{1}$ s.t. $u\equiv 0$ outside a compact set. Then, for a direction $w\in\mathbb{R}^{n}$ s.t. $|w|=1$:
$u(x)=-\int\limits_{0}^{\infty}\partial_{r}u(x+rw)\mathop{}\!\mathrm{d}r.$
The procedure can be seen geometrically as cylinder like coordinates, see Figure 2. By integrating over the unit sphere in $n-k$ dimensions (extended with 00 components in $\mathbb{R}^{n}$) and with $x=0$ and $u(0)=1$ we get:
$c_{n-k}$ $=-\int\limits_{0}^{\infty}\int\limits_{\mathbb{S}^{n-k-1}}\partial_{r}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n-k-1}(w)\mathop{}\!\mathrm{d}r$
$=-\int\limits_{\mathbb{R}^{n-k}}r^{-(n-k-1)}\partial_{r}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n-k}(r\cdot w)$
$=-\int\limits_{\mathbb{R}^{n-k}}\frac{r}{|rw|^{n-k}}w\cdot\nabla^{n-k}u(rw)\mathop{}\!\mathrm{d}\mathcal{H}^{n-k}(r\cdot w)$
$=-\int\limits_{\mathbb{R}^{n-k}}\frac{y\cdot\nabla^{n-k}u(y)}{|y|^{n-k}}\mathop{}\!\mathrm{d}\mathcal{H}^{n-k}(y).$
This allows an estimate for the capacity of partial radially symmetric sets for $u$ as in the definition of $\text{Cap}_{2,\mu}(L_{r},L_{R})$ (Definition 2.6):
11 11 $\simeq\left(\int\limits_{L}\mathop{}\!\mathrm{d}\mathcal{H}^{k}(x)\right)^{2}$
$\simeq\left(\int\limits_{L}\int\limits_{\mathbb{R}^{n-k}(x)}\frac{(x-y)\cdot\nabla^{n-k}u(y)}{|x-y|^{n-k}}\mathop{}\!\mathrm{d}\mathcal{H}^{n-k}(y)\mathop{}\!\mathrm{d}\mathcal{H}^{k}(x)\right)^{2}$
$\lesssim\int\limits_{L_{R}}|\nabla^{n-k}u|^{2}\mathop{}\!\mathrm{d}\mu_{i}\cdot\int\limits_{L_{R}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}(y)}{\omega_{i}(y)}.$
By estimating the gradient pointwise against the lower dimensional gradient, this yields:
$\text{Cap}_{2,\mu_{i}}(L,L_{R})$ $=\text{Cap}_{2,\mu_{i}}(\bigcap\limits_{r>0}L_{r},L_{R})=\lim\limits_{r\searrow 0}\text{Cap}_{2,\mu_{i}}(L_{r},L_{R})$
$\geq\lim\limits_{r\searrow 0}\inf\left\{\int\limits_{L_{R}}|\nabla^{n-k}u|^{2}\mathop{}\!\mathrm{d}\mu_{i}:u\geq\mathds{1}_{L_{r}}\right\}$
$\gtrsim\left(\int\limits_{L_{R}}\frac{1}{\operatorname*{dist}(y,L)^{2(n-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}(y)}{\omega_{i}(y)}\right)^{-1}.$
[…]
Figure 2: We prove the lower capacity bound via a foliation procedure.
Next, we note that in our setting the density $\omega_{2}$ is in such a way that
$\int_{L_{r}}\omega_{2}\mathop{}\!\mathrm{d}\text{vol}_{M_{2}}\int_{L_{r}}\frac{1}{\omega_{2}}\mathop{}\!\mathrm{d}\text{vol}_{M_{2}}\simeq r^{2(n_{2}-k)}$
as it is bounded from above and below. This can be compared to a Muckenhoupt condition on the level of $L$.
The lower bound can now be simplified into a more suitable version and connected to the other manifold. By the layer cake formula and the transformation $\rho=t^{-\frac{1}{2(n-k-1)}}$:
$\int\limits_{L_{R}}$ $\frac{1}{\operatorname*{dist}(y,L)^{2(n-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n}}{\omega}$
$=\int\limits_{0}^{\infty}\frac{\mathcal{L}^{n}}{\omega}(x\in L_{R}:\operatorname*{dist}(x,L)^{-2(n-k-1)}>t)\mathop{}\!\mathrm{d}t$
$=R^{-2(n-k-1)}\frac{\mathcal{L}^{n}}{\omega}(L_{R})+\int\limits_{R^{-2(n-k-1)}}^{\infty}\frac{\mathcal{L}^{n}}{\omega}(L_{t^{-\frac{1}{2(n-k-1)}}})\mathop{}\!\mathrm{d}t$
$=R^{-2(n-k-1)}\frac{\mathcal{L}^{n}}{\omega}(L_{R})+2(n-k-1)\int\limits_{0}^{R}\rho^{-(2n-2k-1)}\frac{\mathcal{L}^{n}}{\omega}(L_{\rho})\mathop{}\!\mathrm{d}\rho.$
To combine the estimates on both manifolds, we now want to compare $\frac{\mathcal{L}^{n}}{\omega}(L_{R})$ (and $\omega\mathcal{L}^{n}(L_{R})$) on both manifolds to get an estimate of $\frac{\mathcal{L}^{n_{1}}}{\omega_{1}}(L_{R})$ by $\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{R})$. To this end, consider the Vitali covering theorem with the covering of $L_{R}$ by $B_{R}(x)$ for $x\in L$. There exist finitely many $x_{j}\in L$ s.t. $L_{R}\subset\bigcup 5B_{j}$ ($5B_{j}:=B_{5R}(x_{j})$) and $B_{R}(x_{j})$ are disjoint. Since $\omega$ is a 2-Muckenhoupt weight, we know that $\omega\mathcal{L}^{n}(B_{R})$ and $\frac{\mathcal{L}^{n}}{\omega}(B_{R})$ can be compared: $\omega\mathcal{L}^{n}(B_{R})\cdot\frac{\mathcal{L}^{n}}{\omega}(B_{R})\simeq R^{2n}$.
Hence,
$\frac{\mathcal{L}^{n_{1}}}{\omega_{1}}(L_{R})$ $\leq\sum\limits_{j}\frac{\mathcal{L}^{n_{1}}}{\omega_{1}}(5B_{j})\lesssim\sum\limits_{j}\frac{\mathcal{L}^{n_{1}}}{\omega_{1}}(B_{j})\lesssim R^{2n_{1}}\sum\limits_{j}(\omega_{1}\mathcal{L}^{n_{1}}(B_{j}))^{-1}$
$\leq N(R)R^{2n_{1}}\sum\limits_{j}(\omega_{2}\mathcal{L}^{n_{2}}(B_{j}))^{-1}\lesssim N(R)R^{2n_{1}-2n_{2}}\sum\limits_{j}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(B_{j})$
$\leq N(R)R^{2n_{1}-2n_{2}}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{R}).$
In these estimates, we used the doubling condition on the weighted manifold $M_{1}$.
Together, this yields:
$\text{Cap}_{2,\mu_{1}}(L,L_{R})$
$\gtrsim\left(R^{-2(n_{1}-k-1)}\frac{\mathcal{L}^{n_{1}}}{\omega_{1}}(L_{R})+2(n_{1}-k-1)\int\limits_{0}^{R}\rho^{-(2n_{1}-2k-1)}\frac{\mathcal{L}^{n_{1}}}{\omega_{1}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}$
$\gtrsim\left(N(R)R^{-2(n_{2}-k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{R})+c\int\limits_{0}^{R}N(\rho)\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}$
$\gtrsim\left(c_{R}+c\int\limits_{0}^{R}N(\rho)\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
With our generalized Muckenhoupt condition on $\omega_{2}$, this has a more direct form and we can close the chain of arguments directly:
$\text{Cap}_{2,\mu_{1}}(L,L_{R})$ $\gtrsim\left(N(R)R^{2}\mu_{2}(L_{R})^{-1}+c\int\limits_{0}^{R}N(\rho)\rho\mu_{2}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1},$
$\text{Cap}_{2,\mu_{2}}(L,L_{R})$ $\gtrsim\left(R^{2}\mu_{2}(L_{R})^{-1}+2(n_{2}-k-1)\int\limits_{0}^{R}\rho\mu_{2}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Since we are in the case where $\omega_{2}$ is bounded and $k=n_{2}-1$, we see that the relevant condition on $N(R)$ is
$\int\limits_{0}^{R}N(\rho)\mathop{}\!\mathrm{d}\rho<\infty$
as we fixed $R>0$ and $\mu_{2}(L_{\rho})\simeq\rho$. The dimensional condition on $k$ is necessary for $\mu_{2}(L)=0$ and $\text{Cap}_{2,\mu_{2}}(L,L_{R})>0$.
Thus, the intersection has positive capacity in $M_{1}$ in this setting if the same holds for $M_{2}$ and $\mu$ satisfies a $N$-doubling condition with integrable $N$.
Since the global and restricted capacity are equivalent for fixed $R>0$, this implies that q.e. point of the intersection is a Lebesgue point, c.f. [40, Theorem 3.5] and [41, Corollary 3.7]. Therefore, one has a trace in the sense of a quasi-continuous refinement on $L$. ∎
3.3 Compact embedding of $H^{1}_{*}$
[…]
We prove the compactness of the embedding through sequentially compactness and an inner approximation up to the boundary. The idea of this proof is based on the techniques of [55, Chapter 5]. For the convenience of the reader, we explain the details in our case here.
Lemma 3.2 (Compact embedding of $H^{1}_{*}$).
Let $(X,d,\mu)$ be as in Theorem 2.2 and $\operatorname*{\mathcal{E}}$ as before. Then, $H^{1}_{*}$ is compactly embedded into $L^{2}(X,\mu)$, i.e. for $(u_{n})_{n\in\mathbb{N}}\subset H^{1}_{*}$ bounded, there exists a subsequence that converges strongly in $L^{2}(X,\mu)$.
Proof.
First, we prove the compactness in the inner of the domain and then extend this result to the whole space using the smoothness of the boundary.
Take $(u_{n})_{n\in\mathbb{N}}\subseteq D(\operatorname*{\mathcal{E}})$ with
$\int\limits_{X}|u_{n}|^{2}\mathop{}\!\mathrm{d}\mu+\operatorname*{\mathcal{E}}(u_{n})\leq C<\infty$
and define $X_{\varepsilon}:=\{x\in X:\operatorname*{dist}(x,\partial X)<\varepsilon\}$ where $\partial X:=\bigcup\partial M_{i}$. We will prove that this sequence has a converging subsequence in $L^{2}$. Choose a open covering of $X\setminus X_{\varepsilon}$ by balls, via doing this for both $M_{i}$ and taking the union. This covering consists of $q$ balls of fixed radius $r<\varepsilon$ s.t.
$q\lesssim\left(\frac{2(\operatorname*{diam}(X)+r)}{r}\right)^{\nu}$
and each point is contained in at most $l\lesssim 2^{4\nu}$ balls where $\nu$ depends on the doubling constants of $\text{vol}_{M_{i}}$ and is independent of $r$. In the following, we call this balls $B_{i}$ where only the balls restricted to the manifolds are meant. They do not need to be metric balls in $X$.
By Banach-Alaoglu there exists a subsequence and $u\in L^{2}$ s.t. $u_{n}\rightharpoonup u$ in $L^{2}$. Let $\omega_{n,m}:=u_{n}-u_{m}$ be the difference. We aim to show $\omega_{n,m}\to 0$ in $L^{2}(X\setminus X_{\varepsilon})$.
With $(\omega_{n,m})_{i}$ as the mean value, we can estimate
$\int\limits_{X\setminus X_{\varepsilon}}\omega_{n,m}^{2}\mathop{}\!\mathrm{d}\mu$ $\leq\sum\limits_{i}\int\limits_{B_{i}}\omega_{n,m}^{2}\mathop{}\!\mathrm{d}\mu$
$\lesssim\sum\limits_{i}\int\limits_{B_{i}}|\omega_{n,m}-(\omega_{n,m})_{i}|^{2}\mathop{}\!\mathrm{d}\mu+\sum\limits_{i}\int\limits_{B_{i}}|(\omega_{n,m})_{i}|^{2}\mathop{}\!\mathrm{d}\mu$
[…]
The first part I can be estimated by the Poincare inequality on balls on $M_{i}$, see Section 2:
$\text{I}=\sum\limits_{i}\int\limits_{B_{i}}|\omega_{n,m}-(\omega_{n,m})_{i}|^{2}\mathop{}\!\mathrm{d}\mu$ $\leq r^{2}\sum\limits_{i}\int\limits_{B_{i}}|\nabla\omega_{n,m}|^{2}\mathop{}\!\mathrm{d}\mu$
$\leq lr^{2}\int\limits_{X}|\nabla\omega_{n,m}|^{2}\mathop{}\!\mathrm{d}\mu$
$\lesssim lr^{2}.$
The second part II directly yields:
II $=\sum\limits_{i}\frac{1}{\mu(B_{i})}\left(\int\limits_{B_{i}}\omega_{n,m}\mathop{}\!\mathrm{d}\mu\right)^{2}$
$\leq q\max\limits_{i}\frac{1}{\mu(B_{i})}\left(\int\limits_{B_{i}}\omega_{n,m}\mathop{}\!\mathrm{d}\mu\right)^{2}$
$\lesssim q\frac{1}{r^{\nu}\mu(X)}\max\limits_{i}\left(\int\limits_{B_{i}}\omega_{n,m}\mathop{}\!\mathrm{d}\mu\right)^{2}.$
Here, we used the doubling condition on $M_{i}$ to compare $\mu(B_{i})$ to $\mu(X)$.
Since $\omega_{n,m}\rightharpoonup 0$ in $L^{2}$ as $n,m\to\infty$ we have $\int\limits_{B_{i}}\omega_{n,m}\mathop{}\!\mathrm{d}\mu\to 0$. Now, fix $\delta>0$ and plug in $r=\sqrt{\frac{\delta}{l}}$:
$\int\limits_{X\setminus X_{\varepsilon}}\omega_{n,m}^{2}\mathop{}\!\mathrm{d}\mu\lesssim\delta+\frac{ql^{\frac{\nu}{2}}}{\delta^{\frac{\nu}{2}}\mu(X)}\max\limits_{i}\left(\int\limits_{B_{i}}\omega_{n,m}\mathop{}\!\mathrm{d}\mu\right)^{2}<2\delta.$
This holds for $n,m$ large enough since $q$ is bounded. Thus, we can find $n_{*},m_{*}$ dependent on $\delta$ s.t. for $n>n_{*},m>m_{*}$ this inequality holds. We conclude that $\omega_{n,m}$ converges to 00 in $L^{2}$.
To approximate $X$, we need that the derivative cannot concentrate on the boundary. We define a quantity that measures the mass which can concentrate on the boundary
$\Gamma_{X}(\varepsilon)$ $:=\sup\limits_{\left\lVert u\right\rVert_{H^{1}_{*}(X)}=1}\int\limits_{X_{\varepsilon}}|u|^{2}\mathop{}\!\mathrm{d}\mu,$
$\Gamma_{X}(0)$ $:=\lim\limits_{\varepsilon\searrow 0}\Gamma_{X}(\varepsilon)$
Since the manifolds are smooth with smooth boundary, we have $\Gamma_{X}(0)=0$.
Take $u_{n}$ a sequence in $L^{2}(X)$ s.t. $\left\lVert u_{n}\right\rVert_{H^{1}_{*}(X)}\leq 1\;\forall n$. Then, for every $\delta>0$ we find $\varepsilon$ s.t. $\Gamma_{X}(\varepsilon)<\delta$ and a subsequence $u_{n}$ s.t.
$u_{n}$ $\rightharpoonup u\text{ in }L^{2}(X),$
$\forall n\geq n_{0}(\varepsilon)$ $:\left\lVert u-u_{n}\right\rVert_{L^{2}(X\setminus X_{\varepsilon})}^{2}\leq\delta.$
The sequence is constructed via a diagonal sequence. First, take a subsequence $u_{n}\rightharpoonup u$ in $L^{2}(X)$. Afterwards, for a sequence $\varepsilon\to 0$ we can find refining subsequences (the existence is given by a diagonal sequence) s.t. $u_{n}\to u$ in $L^{2}(X\setminus X_{\varepsilon})$ and for $n\geq n_{0}(\varepsilon)$:
$\left\lVert u-u_{n}\right\rVert_{L^{2}(X\setminus X_{\varepsilon})}^{2}\leq\delta.$
Next, we can compute the norm to show strong convergence in $L^{2}(X)$:
$\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X)}^{2}$ $=\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X\setminus X_{\varepsilon})}^{2}+\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X_{\varepsilon})}^{2}$
$\lesssim\delta+\limsup_{n\to 0}\left\lVert u_{n}\right\rVert_{L^{2}(X_{\varepsilon})}^{2}+\left\lVert u\right\rVert_{L^{2}(X_{\varepsilon})}^{2}$
$\lesssim\delta+\limsup_{n\to 0}\left\lVert u_{n}\right\rVert_{L^{2}(X_{\varepsilon})}^{2}$
$\lesssim\delta+\delta\left\lVert u_{n}\right\rVert_{H^{1}_{*}(X)}^{2}\lesssim\delta$
We used the lower semi-continuity of the $L^{2}$ norm w.r.t. to the weak $L^{2}$ convergence. With $\delta\to 0$ this implies the compact embedding. ∎
Therefore, the image of the resolvent $R_{\operatorname*{\mathcal{A}}}$ is compactly embedded into $L^{2}$ which yields that $R_{\operatorname*{\mathcal{A}}}$ is a compact resolvent.
Now, we can combine these calculations and conclude that only constants lie in the kernel of $A$.
Given the assumptions above, $H^{1}_{*}$ has a continuity condition. Thus, assume $f\in\text{Ker}(A)$. Then
$0=-(Af,f)=\operatorname*{\mathcal{E}}(f).$
On each manifold, we fix charts and take partition of unities on the charts which by the transformation to $\mathbb{R}^{n}$ yield
$\left\lVert\nabla f\right\rVert_{L^{2}(\mu)}=0.$
Since $\omega_{i}>0$, we conclude that $f$ is constant a.e. on each chart. By approximation and since the manifolds are connected, $f$ is constant on each manifold. Now, we can infer the value of $f$ on $L$ q.e. by the mean-value on each manifold and it has to be the same on both sides (since the capacities are equivalent), which yields that $f$ has to be constant on $X$, i.e. the kernel of $A$ are the constant functions. Together with the facts that $A$ is a selfadjoint, non-positive operator and has a compact resolvent by operator-theory, it follows that the spectrum of $A$ is discrete with no finite accumulation point and $A$ has a spectral gap, c.f. [64, Chapter 8]. This is equivalent to satisfying a global Poincare type inequality on the orthogonal complement of the constants $\langle 1\rangle^{\perp}$. To see this, we can write down the condition on the orthogonal complement of the kernel. There exists $\lambda\geq 0$ s.t.
$\lambda\left\lVert f-f_{X}\right\rVert_{L^{2}(X,\mu)}\leq-(Af,f)=\mathcal{E}(f).$
Next, we apply the spectral theorem which represents $A$ as a multiplication operator $a\leq 0$ on $L^{2}$ with $a\leq-\lambda$ on the orthogonal complement of the constants. For f with $\int f\mathop{}\!\mathrm{d}\mu=0$ this yields
$\left\lVert S_{h}f\right\rVert_{L^{2}(\mu)}\leq e^{-\lambda h}\left\lVert f_{0}\right\rVert_{L^{2}(\mu)}\to 0.$
Therefore, solutions converge exponentially fast to the constants that are their mean value.
This, in turn, proves the irreducibility or ergodicity of $S_{h}$ following the Definitions 2.7 and 2.8.
Proof of Theorem 2.2.
First, we prove the ergodicity. If $E$ satisfies $T_{t}^{-1}(E)=E$ then we also have
$T_{t}(E)=T_{t}(T_{t}^{-1}(E))\subseteq E.$
By monotonicity we know that if $A\subseteq B$ then
$T_{t}(A)\subseteq T_{t}(B).$
This holds as $T_{t}$ is monotone since $S_{h}$ is linear and monotone (positive). By [20, Theorem 1.4.1], $S_{h}$ is Markovian as it is the semigroup associated to a Dirichlet form, i.e. if $0\leq f\leq 1$ then $0\leq S_{h}\leq 1$. Thus, by linearity of $S_{h}$ it is positivity preserving on a.e. bounded functions and therefore monotone.
By the upper bound ($S_{t}\mathds{1}_{A}\leq 1$ a.e.), $S_{s}\mathds{1}_{A}\leq T_{s}(A)$ and thus
$T_{t+s}(A)\subseteq T_{t}\circ T_{s}(A).$
If $\mu(E)>0$ we have $T_{t}(E)\to X$ as $t\nearrow\infty$ by the observation that all functions converge to their mean as a constant function under the flow. If $T_{t}(E)\subseteq E$ for $t>0$ then
$T_{2t}(E)\subseteq T_{t}\circ T_{t}(E)\subseteq T_{t}(E)\subseteq E.$
Inductively this implies that a.e. point of $X$ is contained in $E$ and thus $\mu(E^{c})=0$.
To prove irreducibility, we can plug in $f\equiv 1$ and get $S_{h}(\mathds{1}_{E})=\mathds{1}_{E}$ $\mu-$a.e. which implies $T_{h}(E)=E$ a.e.. ∎
4 Generalizations
In this section, we show how the proof and techniques of Section 3 can be generalized to the case of more then two connected manifolds. Moreover, we give an idea how to incorporate more general weights.
We still rely on the equivalence of the capacities to get a trace in the capacity of our Sobolev space. Thus, the effective dimension (taking into account the weight) of the manifolds at their intersection cannot vary too much. Heuristically spoken, the difference of $k$ and the effective dimension minus one has to be less then one.
For general weights $\omega_{1},\omega_{2}\in A_{2}$, we can follow the same calculations to get lower bounds on the capacities:
$\text{Cap}_{2,\mu_{1}}(L,L_{R})$ $\gtrsim\left(N(R)R^{-2(n_{2}-k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{R})\right.$
$\quad\left.+2(n_{1}-k-1)\int\limits_{0}^{R}N(\rho)\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1},$
$\text{Cap}_{2,\mu_{2}}(L,L_{R})$ $\gtrsim\left(R^{-2(n_{2}-k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{R})\right.$
$\quad\left.+2(n_{2}-k-1)\int\limits_{0}^{R}\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Additionally, if $\omega_{2}$ satisfies the $L-$Muckenhoupt type condition on the intersection,
$\int_{L_{r}}\omega_{2}\mathop{}\!\mathrm{d}\text{vol}_{M_{2}}\int_{L_{r}}\frac{1}{\omega_{2}}\mathop{}\!\mathrm{d}\text{vol}_{M_{2}}\simeq r^{2(n_{2}-k)},$
we can simplify these expressions to
$\text{Cap}_{2,\mu_{1}}(L,L_{R})$ $\gtrsim\left(N(R)R^{2}\mu_{2}(L_{R})^{-1}+2(n_{1}-k-1)\int\limits_{0}^{R}N(\rho)\rho\mu_{2}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1},$
$\text{Cap}_{2,\mu_{2}}(L,L_{R})$ $\gtrsim\left(R^{2}\mu_{2}(L_{R})^{-1}+2(n_{2}-k-1)\int\limits_{0}^{R}\rho\mu_{2}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Example.
We consider a submanifold $L$ in $M$ with $n:=\dim M$ and $k:=\dim L\neq n-1$ with weight $\omega$ which locally around $L$ has the form
$\omega(x):=\frac{1}{\operatorname*{dist}(x,L)^{\alpha}}.$
Then locally,
$\mu_{i}(L_{r})$$\simeq\begin{cases}\frac{r^{n-k-\alpha}}{n-k-\alpha}&\text{if }\alpha<n-k,\\ \infty&\text{else,}\end{cases}$
$\frac{\mathcal{L}^{n}}{\omega}=\int\limits_{L_{r}}\frac{1}{\omega}\mathop{}\!\mathrm{d}x$$\simeq\begin{cases}\frac{r^{n-k+\alpha}}{n-k+\alpha}&\text{if }\alpha>-(n-k),\\ \infty&\text{else.}\end{cases}$
Thus, as $\mu$ needs to be locally finite and the capacity needs to be finite, we require $\alpha\in(-(n-k),n-k)$. In this case, the $L-$Muckenhoupt type condition is satisfied and the requirement for the capacities becomes
$\int\limits_{0}^{R}\rho\mu_{i}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho$ $=\int\limits_{0}^{R}\rho^{1-n+k+\alpha}\mathop{}\!\mathrm{d}\rho<\infty,$
i.e. $\alpha>n-k-2$ which coincides with our intuition that the difference of $k$ and effective dimension minus one, $n-\alpha-1$, has to be less than one.
If $L$ is the intersection with another manifold and the whole measure $\mu$ satisfies the $N-$doubling condition, the condition for the other manifold is
$\int\limits_{0}^{R}N(\rho)\rho^{1-n+k+\alpha}\mathop{}\!\mathrm{d}\rho<\infty.$
Moreover, we can prove an upper capacity bound which shows that our lower bound is optimal up to a constant. This upper bound of the capacity can be obtained through an explicit competitor. First, consider $L_{r}$ in $L_{2r}$ and the linear competitor $u(x)=\left(1-\frac{\operatorname*{dist}(x,L_{r})}{r}\right)\mathds{1}_{L_{2r}}\in\operatorname*{Lip}\cap C_{c}(L_{2r})$:
$\text{Cap}_{2,\mu_{2}}(L_{r},L_{2r})$ $\lesssim\int\limits_{L_{2r}\setminus L_{r}}r^{-2}\mathop{}\!\mathrm{d}\mu_{2}$
$\leq\int\limits_{L_{2r}}r^{-2}\mathop{}\!\mathrm{d}\mu_{i}$
$\lesssim r^{2n_{2}-2k-2}\left(\int\limits_{L_{2r}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{2}}}{\omega_{2}}\right)^{-1}$
$\leq r^{2n_{2}-2k-2}\left(\int\limits_{L_{2r}\setminus L_{r}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{2}}}{\omega_{2}}\right)^{-1}$
$\leq\left(\int\limits_{L_{2r}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{2}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{2}}}{\omega_{i}}\right)^{-1}.$
This can be used iteratively to compute the capacity. With $k\in\mathbb{N}$ s.t. $2^{k-1}r\leq R\leq 2^{k}r$ we get:
$\text{Cap}_{2,\mu_{i}}(L_{r},L_{R})$ $\lesssim\text{Cap}_{2,\mu_{i}}(L_{r},L_{2^{k}}r)$
$\leq\left(\sum\limits_{j=0}^{k-1}\text{Cap}_{2,\mu_{i}}(L_{2^{j}r},L_{2^{j+1}r})^{-1}\right)^{-1}$
$\leq\left(\int\limits_{L_{2^{k}r}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{i}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}}{\omega_{i}}\right)^{-1}$
$\leq\left(\int\limits_{L_{R}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{i}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}}{\omega_{i}}\right)^{-1}.$
The proof of the first inequality is analogous to [36, Lemma 2.16]. With the simplification from Section 3,
$\text{Cap}_{2,\mu_{i}}(L,L_{R})\lesssim\left(R^{-2(n_{i}-k-1)}\frac{\mathcal{L}^{n_{i}}}{\omega_{i}}(L_{R})+2(n_{i}-k-1)\int\limits_{0}^{R}\rho^{-(2n_{i}-2k-1)}\frac{\mathcal{L}^{n_{i}}}{\omega_{i}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
With $\omega_{i}$ satisfying the Muckenhoupt type condition, this yields
$\text{Cap}_{2,\mu_{i}}(L,L_{R})$ $\lesssim\left(R^{2}\mu_{i}(L_{R})^{-1}+2(n_{i}-k-1)\int\limits_{0}^{R}\rho\mu_{i}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Next, we prove the iteration procedure for more than two manifolds. We are in the setting of $(X,d,\mu)$ where $X$ is given as the connected union of embedded Euclidean manifolds of varying dimension glued at smooth submanifolds:
$X:=\bigcup\limits_{i=1}^{m}M_{i}.$
Here, the manifolds $M_{i}$ are compact, connected, oriented with smooth boundary and of dimension $n_{i}$. We require that their pairwise intersections $L_{ij}$ are disjoint to each other, lie in the inner of $M_{i},M_{j}$ and are transversal. Additionally, they have to satisfy $\mu(L_{ij})=0$ and that $X$ is path-wise connected. The transversality does not need to be satisfied in the whole space that $X$ is contained in, a local transversality is suitable, i.e. there exists a neighborhood manifold $Y$ containing $\left(L_{ij}\right)_{\varepsilon}$ for some $\varepsilon>0$ s.t. $M_{i}$ and $M_{j}$ intersect transversal in $Y$. We denote the dimension of $L_{ij}$ by $k_{ij}$.
The measure
$\mu:=\sum\limits_{i=1}^{m}\omega_{i}\cdot\text{vol}_{M_{i}}$
and Dirichlet form
$\operatorname*{\mathcal{E}}(f):=\int\limits_{X}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu:=\sum\limits_{i=1}^{m}\int\limits_{M_{i}}|\nabla^{M_{i}}f|^{2}\mathop{}\!\mathrm{d}\mu_{i}$
are defined analogously to before. In particular, $\omega_{i}\in A_{2}(M_{i}),\omega_{i}>0$. Again, we can choose $d$ to be the glued distance. Here, $\operatorname*{\mathcal{E}}$ is strictly local and regular and $\mu$ has to satisfy a $N-$doubling condition with $N\in L^{1}_{loc}$.
Additionally, assume that there are neighborhoods of $L_{ij}$ s.t. $\mu_{j}$ is bounded from above and below and that $k_{ij}+1=n_{j}<n_{i}$.
In this setting, we can use the same proofs as before as the calculations for the capacities are done locally around the intersections and the other proofs are done analogously for a finite number of manifolds. I.e. we still can use that $(X,d)$ is a metric doubling space to prove the compact embedding and with the connectedness and capacity conditions, the ergodicity and irreducibility follow analogously.
5 Outlook
In this section, we will give the needed notations and definitions to define a non-local perimeter functional, the heat excess $E_{h}$. An interesting question is whether this functional converges to the perimeter as $h$ goes to zero. This proven in the context of Euclidean space (e.g. [19]) and smooth manifolds ([49]) but unknown in our setting or general metric measure spaces. In this context, it is known that the limit leads to an equivalent characterization of BV under assumptions on the measure, c.f. [1, 2, 3, 52, 58, 59].
We define the perimeter functional for a function $f\in L^{1}(X,\mu)$ as (in accordance with [52]):
$\left\lVert Df\right\rVert(U):=\liminf\left\{\left\lVert\nabla f_{n}\right\rVert(U)\left|f_{n}\to f\text{ in }L^{1},f_{n}\in\operatorname*{Lip}\right.\right\}.$
In this note, we use $\nabla f$ if the density exists as a function (absolutely continuous) and $Df$ for a general derivative or in this case the perimeter / total variation.
Additionally, we define the heat excess at time $h>0$ via
$E_{h}(f):=\int\limits_{X}S_{h}(x,y)|f(x)-f(\cdot)|(x)\mathop{}\!\mathrm{d}\mu(x).$
If the heat semigroup $S_{h}$ possesses a heat kernel $p_{h}(x,y)$, we can rewrite it as
$E_{h}(f):=\int\limits_{X}\int\limits_{X}p_{h}(x,y)|f(x)-f(y)|\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
If $f=\chi_{U}$ is a characteristic function, it becomes
$\mathcal{E}_{h}(f)=\int\limits_{X}\chi_{U^{C}}S_{h}\chi_{U}\mathop{}\!\mathrm{d}\mu(x).$
This can also be seen as a Korevaar-Schoen type energy. Thus, the convergence is related to the Bourgain-Brezis-Mironescu formula, c.f. [31].
An interesting question is whether the following conjecture holds true in general or even whether it is true in the setting of glued manifolds with $A_{1}$ weights and doubling measure.
Conjecture (Perimeter functional).
Let $(X,d,\mu)$ be a complete metric space with $(\operatorname*{\mathcal{E}},D(\operatorname*{\mathcal{E}}))$ a strictly local, regular Dirichlet form. Moreover, assume that $\mu$ satisfies a $1-$Poincare inequality and is doubling. Then, for $f\in BV(X)$:
$E_{h}(f)\overset{L^{1}-\Gamma}{\longrightarrow}\left\lVert Df\right\rVert(X).$
Given this convergence, we can construct a classifier by considering a minimizing movement scheme similar to the thresholding scheme (MBO-scheme) to receive a fast iterative algorithm to classify high-dimensional data.
In proving the $\Gamma-$convergence of the heat excess $E_{h}(f)\overset{\Gamma}{\to}\left\lVert Df\right\rVert$ in $L^{1}$, it is necessary for $\mu$ to satisfy a $1-$Poincare inequality (or similar condition on the measure $\mu$).
Definition 5.1 ($p-$Poincare inequality).
We say that $(X,d,\mu)$ satisfies a $(q,p)-$Poincare inequality if, for each $f\in\operatorname*{Lip}$, there exists $C>0$ and a dilation factor $\lambda\geq 1$ s.t. for each ball $B$:
$\left(\fint_{B}|f-f_{B}|^{q}\mathop{}\!\mathrm{d}\mu\right)^{\frac{1}{q}}$ $\leq Cr\left(\fint_{\lambda B}|\nabla f|^{p}\mathop{}\!\mathrm{d}\mu\right)^{\frac{1}{p}}.$
If $p=q$, we will call this condition also a $p-$Poincare inequality.
Remark.
If $X$ satisfies a $(q,p)-$Poincare inequality then the same is true for $1\leq q^{\prime}\leq q$ but we also can recover a $(p,p)-$Poincare inequality in doubling spaces. Moreover, if $X$ satisfies a $(1,p)-$Poincare inequality, it satisfies a $(1,p^{\prime})-$Poincare inequality for each $1\leq p^{\prime}\leq p$.
In the setting of glued manifolds, for a $1-$Poincare inequality to hold on the whole space, it is required to hold on the components and a $1-$capacity condition needs to be satisfied.
Definition 5.2 ($p-$Capacities).
The relative $p-\mu-$capacity is defined on compact sets and extended to measurable sets in the usual way:
$\text{Cap}_{p,\mu}(K,\Omega):=\inf\left\{\left.\int\limits_{\Omega}|\nabla u|^{p}\mathop{}\!\mathrm{d}\mu\right|u\in\operatorname*{Lip}\cap C_{c}(\Omega),u\geq\mathds{1}_{K}\right\}.$
Moreover, with $1-$Muckenhoupt weights, the $1-$Poincare inequality is satisfied on the individual manifolds.
Definition 5.3 ($p-$Muckenhoupt weights).
We say $\omega$ is in the class $A_{p}(M)$ of $p-$Muckenhoupt weights on a smooth compact manifold $M$ for $p>1$ if for all balls $B$
$\fint\limits_{B}\omega_{i}\mathop{}\!\mathrm{d}\text{vol}_{M_{i}}\cdot\left(\fint\limits_{B}\frac{1}{\omega_{i}^{\frac{1}{p-1}}}\mathop{}\!\mathrm{d}\text{vol}_{M_{i}}\right)^{p-1}\lesssim 1$
and $\omega\in A_{1}(M)$ if
$\fint\limits_{B}\omega_{i}\mathop{}\!\mathrm{d}\text{vol}_{M_{i}}\simeq\operatorname*{ess\,inf}_{B}\omega.$
For the future, an aim is to investigate the $\Gamma-$ or Mosco–convergence of $E_{h}$ to $\left\lVert Df\right\rVert$. To this end, different inequalities as well as bounds for the heat kernel are interesting. It is notable that one has to consider $A_{1}$ weights and $1-$Poincare inequalities instead of the $L^{2}$ variants for this. Moreover, to carry over these inequalities from the components to the whole space, one needs more refined geometric assumptions on the intersection related to $\text{Cap}_{1,\mu}$.
Another aim is to investigate more refined properties of the heat flow and associated Brownian motion. Additionally, for the iteration and general gluing, even at already existing intersections, it is interesting to investigate the assumptions on the glued space, i.e. whether the inequalities like the 2-Poincare inequality are still satisfied globally.
6 Related work
In this work, we consider the setting of glued manifolds with weights. This setting originates from manifold learning and examples of similar approaches were studied before in the works we list next. In [15, 54] the space $\mathbb{R}^{2}\times\mathbb{R}_{+}$ with a collapsed center is analyzed and an expansion of the heat kernel and Brownian motion is computed. Similarly, [50, 51] study so-called distorted Brownian motion in $\mathbb{R}^{3}\times\mathbb{R}_{+}$ where the connection is done via a density that fulfills our conditions. Using the same setting, in [54], Ooi extends the results of [15] to the case of two arbitrary $\mathbb{R}^{d_{1}},\mathbb{R}^{d_{2}}$ meeting at one collapsed point with explicit heat kernel bounds. The works [11, 16] study bow ties as special structures and how Poincare inequalities on them can be obtained from capacity conditions. In [55], glued spaces with constant dimension are analyzed. A major difference is that only manifolds of the same dimension are considered and they don’t posses a density. Finally, in [28], the authors looks at heat kernels on manifolds with ends.
The theory of Muckenhoupt weights and related study of them can be found in a various papers and books. The main sources for this work were [40], [36, Chapter 2 and 15] and [61, Chapter 5] where they are introduced and studied together with the related capacities. For the setting of metric measure spaces, one can see [46] or in the context of non-doubling measures [42].
In their series of four paper [6, 7, 5, 8] the authors investigate, among other things, Muckenhoupt weights on manifolds, where the appendix in [7] is especially noteworthy.
Related to our construction is the theory of Dirichlet forms and Markov processes in [20] and linear operators in [38] as well as capacities, see [40, 41].
The theory of heat flows and kernels with their estimates is extensive and contained in a lot of fields. In [26], Grigor’yan and co-authors provide a condition for the reverse doubling condition on manifolds and prove connections between measure doubling conditions and heat kernel estimates. In [21], he provides a general theory of upper and lower bounds of heat kernels and how to achieve these in a general setting. Other estimates on heat kernels by Grigor’yan are proven in [22, 23, 24, 25, 27].
In [63] and [62], Sturm shows heat kernel bounds under a doubling assumption on the measure. For the bounds on heat kernels, [14] provides theorems which imply an upper bound given a Sobolev inequality of Nash type. This Nash inequality on the manifolds can be proven using [45] where a proof is given for weighted manifolds which fit our setting. Other Nash inequalities can be found in [9, 39].
In [44], a heat flow and Brownian motion are constructed from Dirichlet forms which arise as limits of discrete forms. For the expansion of the heat kernel on manifolds with smooth weights, a parametrix construction is needed which is done in [57]. Even in a more general setting, the heat kernel localizes which can be seen in locality theorems and a comparison result in [56]. In [10], the Lipschitz continuity of heat kernels in metric measure spaces is investigated. More bounds on the heat kernel can be found in [12, 17].
In [33], Hajłasz defines different notions of Sobolev spaces and in [34], he proves for instance Sobolev embeddings of them. A short summary of Hajłasz Sobolev-spaces is given in [32] and [32]. Generalizations of embeddings of Sobolev spaces can be found in [30, 37]. Here, the former analyzes the setting of embeddings in weighted Sobolev spaces whereas the latter gives a generalization of the Rellich-Kondrachov embedding for the Hajłasz Sobolev-space. This proof uses a similar criteria to our $N$-doubling condition.
Functions of bounded variations possess numerous different characterizations in metric spaces. The works [53] and [13] give an overview and characterizations of different definitions. Moreover, in [4] a BMO characterization of sets of finite perimeter is given.
The convergence of the heat excess was already studied in different cases. The most relevant ones to this setting are [1, 2, 3, 19, 49, 52, 58, 59]. First, in [19] the convergence in Euclidean space for the multi-phase mean curvature flow setting is shown. In [49], the $\Gamma$-convergence was proven in the setting of smooth manifolds. [1] as well as [2], [3] and [59] use a weak Bakry Emery condition in the setting of metric measure spaces with a $2$-Poincare inequality to establish equivalence of the limit with the BV seminorm. In comparison, [52] and [58] prove similar results using a $1$-Poincare inequality. The latter is in the setting of smooth manifolds where pointwise convergence is shown.
Similar to this question, is the one after a Bourgain-Brezis-Mironescu formula in metric measure spaces, i.e. whether a non-local kernel perimeter converges to the perimeter. This was investigated in [18, 31, 35, 43, 47, 48].
7 Acknowledgement
I would like to thank my supervisor Prof. Felix Otto for many fruitful discussions and help in this work. Many thanks also to Prof. Tim Laux for the discussions we had and his support.